3. You own a firm that produces sail boats. Sail boats sell for $50,000. You take out a $1 MM loan at a nominal interest rate of 7%, repayable in a year, which will allow you to produce 25 sail boats.

(a) Assume 0% inflation. What is your profit after repayment of the loan?

(b) Assume 3% inflation. What is your profit after repayment of the loan? What real interest rate did you face?

(c) Assume 3% deflation. What is your profit after repayment of the loan? What real interest rate did you face?
Answer #1

(3-) Loan amount o Principal amount $1000000 Nominal interest gati (i) = 7% Number 04, boat produced using the loan amount ((6) Price of boat 50000 + 3% inflation 50000 + 3% x 50000 = 50000 + 0.03 x 500.00 = 50000 + 1500 = $ 51500 Total Revenue from( Inflation rate - 3% Price of boat = S0000+ (-3%) inflation = 50000 __3%x. 50000 = 50000 0:03 X 50000 = S0000 Isoo = $48.500